The width of a rectangle is 3ft less than the length. The perimeter is 10 feet. Find the L & W.
Answer by richwmiller(17219)

The width of a rectangle is 3ft less than the length. The perimeter is 10 feet. Find the L & W.
Let L=length and W=width
W=L-3
L+L+W+W=10
L+L+(L-3) +(L-3)=10
4L-6=10
and 6 to both sides
4L=16
L=4: Length =4
W=4-3
W=1:Width=1
check
4+4+1+1=10
10=10